### v3

# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

Abstract
In this paper, we present a novel framework for enhancing the performance of Quantum Convolution Neural Networks, a.k.a Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and addressing the critical challenges associated with them. Traditional quanvolutional layers, although beneficial for feature extraction, have largely been static, offering limited adaptability. Unlike state-of-the-art, our research overcomes this limitation by enabling training within these layers, significantly increasing the scaling and learning potential of QuNNs. However, in deep QuNNs, i.e., QuNNs with multiple trainable quanvolutional layers induce complexities in gradient-based optimization, primarily due to the restricted gradient flow across these layers. To resolve this, we propose a novel architecture, Residual Quanvolutional Neural Networks (ResQuNNs), leveraging the concept of residual learning, which facilitates the flow of gradients by adding skip connections between these layers. The insertion of residual blocks between quanvolutional layers ensures enhanced gradient access throughout the network, thus enabling deep learning in QuNNs and eventually leading to improved training performance. Moreover, we provide empirical evidence on the strategic placement of these residual blocks within ResQuNNs. Through extensive experimentation, we identify efficient configurations of residual connections, which enable gradients flow across all the layers in the network that eventually results in efficient training. Our results mark a substantial step forward in the evolution of quantum deep learning, offering new avenues for both theoretical development and practical quantum computing applications.
I Introduction
Quantum Convolutional Neural Networks (QCNNs) represent an innovative fusion of quantum computing principles with traditional convolutional neural network (CNN) architectures [1, 2]. These hybrid classical-quantum models utilize parameterized quantum circuits (PQCs) a.k.a ansatz, composed of parameterized rotation gates such as $R_{x}$, $R_{y}$, or $R_{z}$, functioning analogously to convolutional and pooling operations in classical CNNs [3, 4]. The optimization, however, is performed on classical computers [5, 6].
A notable advantage of QCNNs lies in their ability to harness the vast Hilbert space offered by quantum mechanics, which surpasses the capabilities of classical CNNs [7]. This feature allows QCNNs to capture spatial relationships in image data more effectively. Furthermore, the potential future inclusion of higher-level quantum systems, such as qutrits or ququads [8, 9], promises even more sophisticated image comprehension capabilities.
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Fig. 1, focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology.
QuNNs have widely been used in the literature for various applications such as 2D and 3D radiological image classification [12], analysing the high energy physics data [13], for object detection and classification [14, 15]. In light of the capabilities and constraints of present-day quantum hardware, this paper also delves into QuNNs. Our primary focus is on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
Figure 1: A typical Illustration of Quanvolutional Neural Network Architecture.
Figure 2: Motivational analysis showing the impact Trainable quanvolutional layers. When the quantunvolutional layer is untrainable (UT) and only the classical linear layer is trained (T), the QuNN yields suboptimal performance. When the quanvolutional layer is made trainable, it significantly enhanced the training performance. Quanv = Quanvolutional layer, Linear = fully connected layer at the end.
I-A Motivational Analysis
QuNNs, as shown in Fig. 1 performs the convolution on the input image within the grey-shaded region. This process involves sequentially encoding a small region of the input image into quantum space, while moving a kernel across the image, indicated by the green-shaded area of Fig. 1. A unitary transformation is then applied to a small square region of the input via the PQC a.k.a ansatz, as depicted in the brown region of Fig. 1. The output of this quanvolution (a quantum circuit-based convolution) is then processed by a classical neural network to produce the final network output. While this architecture is standard in many state-of-the-art works, a key challenge is the untrainable nature of the quanvolution operation, primarily used for feature extraction. Consequently, the learning process is primarily driven by the subsequent classical network. For QuNNs to be fully scalable, ensuring the trainability of underlying quanvolutional layers is crucial.
In Fig. 2, we assess the impact of trainable versus untrainable quanvolutional layers in QuNNs. To this end, we train QuNNs with both trainable (allowing weight updates during backpropagation) and untrainable (frozen weights, no updates) quanvolutional layers. This training is performed on $1000$ randomly chosen images from MNIST dataset over $30$ iterations, using an $80\%-20\%$ training-validation split. The Adam optimizer with a $0.01$ learning rate was employed. We utilized a $2\times 2$ kernel resulting in $4$-qubit quanvolutional layers, with each layer comprising of $4$ parameterized $R_{y}$ gates and $4$ $CNOT$ entanglement gates. The depth of quantum circuit in the quanvolutional layer is set to $5$.
Results showed suboptimal performance when quanvolutional layers were untrainable and only the final fully connected (FC) layer was trained. Conversely, allowing weight updates in the quanvolutional layers significantly improved learning, enhancing model accuracy by approximately $36\%$. This empirical evidence robustly substantiates the hypothesis that facilitating the trainability of quanvolutional layers significantly augments the model’s overall learning efficacy.
Figure 3: Gradients Propagation in QuNNs for Two (upper pannel) and Three (lower panel) Quanvolutional Layers.
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Fig. 2, scalability of QuNNs requires multiple trainable layers. Also, the rationale for using multiple layers is to avoid the problem of vanishing gradients a.k.a Barren Plateaus (BP), common in single, more expressive PQCs with a higher qubit count and sufficient depth[16, 17, 18]. Multiple layers, with progressively smaller quantum circuits, can potentially mitigate the BPs[19]. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Fig. 3 demonstrates this by showing gradients for two and three cascaded trainable quanvolutional layers.
We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the final layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge in the pursuit of scalable and effective QuNN architectures.
Figure 4: Overview of our contributions
Figure 5: Detailed Methodology. A comprehensive overview of steps involved in the analysis of trainable QuNNs and proposed ResQuNNs. The results of quanvolutional layers are postprocess using both classical layer and quantum circuit. Zero Padding is performed to match the dimensions of input and output before passing it to residual block. The gradients accesibility through all the layers and trainaing and validation accuracy are used as eveluation metrics. QCL = Quanvolutional Layer
I-B Novel Contributions
Our primary contributions encompass the introduction of trainable quanvolutional layers, identification of challenges with multiple quanvolutional layers, the development of Residual Quanvolutional Neural Networks (ResQuNNs) to address these challenges, and empirical findings on the strategic placement of residual blocks to further enhance training performance. An overview of our contributions is shown in Fig. 4.
  • Trainable Quanvolutional Layers: We introduce the concept of making quanvolutional layers trainable in QuNNs, addressing the challenge of enhancing their adaptability within QuNNs and improving their training performance.
  • Challenges with Multiple Quanvolutional Layers: We identify a key challenge in optimizing multiple quanvolutional layers, highlighting the complexities associated with accessing gradients across these layers. In QuNN architectures with multiple trainable quanvolutional layers, the gradients do not flow through all the layers of the network, making deeper QuNNs unsuitable for gradient-based optimization techniques.
  • Residual Quantum Convolutional Neural Networks (ResQuNN): To overcome the gradient accessibility issue, this paper introduces Residual Quantum Convolutional Neural Networks (ResQuNN), drawing inspiration from classical residual neural networks.
  • Residual Blocks for Gradient Accessibility: In ResQuNNs, the residual blocks between quanvolutional layers are incorporated to facilitate comprehensive gradient access throughout the network, thereby enabling the deep learning in QuNNs and eventually improving training performance of QuNNs. These residual blocks combine the output of existing quanvolutional layers with their own or previous layer’s input before forwarding it to the subsequent quanvolutional layer.
  • Strategic Placement of Residual Blocks: We conducted extensive experiments to determine appropriate locations for inserting residual blocks in ResQuNNs which ensure comprehensive gradient throughout the network .
II ResQuNN Methodology
The detailed overview of ResQuNN’s methodology is presented in Fig. 5. Below we discuss different steps of our methodology in detail.
II-A Trainable Quanvolutional Layers in QuNNs.
The first key aspect of our methodology involves the concept of making quanvolutional layers trainable. The majority of works in the literature utilizes these layers for efficient feature extraction and pass them to classical layers, however, the quanvolution operation itself is not trainable. We recognize the need to enhance the adaptability of these layers within QuNNs and introduce trainable quanvolutional layers to harness their full potential. Making the quanvolutional layer trainable entails making the quantum circuit (labeled as 1 in Fig. 5) trainable. We do this by making the quanvolution layer as a pytorch layer which is later used as a standard layer during the model construction. The trainable QuNN utilized in this study is depicted in purple shaded region of Fig. 5, and operates through the following steps.
  1. 1.
    Encoding Classical Data into Quantum States. We use a kernel of size $2\times 2$ which slides over the entire input image with stride 2. This results in $4$ classical features extracted after every step, i.e., $x=\{x_{1},x_{2},x_{3},x_{4}\}$. These features are then encoded into quantum states, specifically into the rotation angles of $RY$ gates. The initial state of each qubit |$\ket{q_{i}}$ is the ground state |$\ket{0}$, and the RY gate applies a rotation by $x_{i}$ radians on the Bloch sphere, which can be described by the following Equation:
    $U_{{\text{Enc}}_{1}}=|q_{i}\rangle=RY(x_{i})|0\rangle=\cos\left(\frac{x_{i}}{2}\right)|0\rangle+\sin\left(\frac{x_{i}}{2}\right)|1\rangle$ (1)
    A total of $4$ qubits will be required, for encoding $4$ input features:
    $U_{{\text{Enc}}_{1}}=\bigotimes_{i=1}^{4}RY(x_{i})\ket{0}$ (2)
  2. 2.
    Trainable Quantum Circuit Operations. The encoded quantum states are then subjected to unitary evolution via an $RX$ rotation gate. Let $\theta_{i}$ be the angle of rotation for the $i$-th qubit applied by the $RX$ gate:
    $U_{\text{Rotation}}=RX(\theta_{i})|q_{i}\rangle=\cos\left(\frac{\theta_{i}}{2}\right)|q_{i}\rangle-i\sin\left(\frac{\theta_{i}}{2}\right)|q_{i}\rangle$ (3)
    For $4$-qubit layer, the above equation can written as:
    $U_{\text{Rotation}}=\bigotimes_{i=1}^{4}RX(\theta_{i})$ (4)
    To introduce entanglement between qubits in the quanvolutional layers, CNOT gates are applied between nearest neighbor qubits. The entanglement operation in this context can be defined as:
    $U_{\text{{Entangle}}}=\bigotimes_{j=1}^{3}CNOT_{j,j+1}$ (5)
  3. 3.
    Overall Transformation for Quantum Circuit. Combining the above steps, the overall transformation $U_{\text{Circuit}_{1}}$ applied by the quanvolutional layer can be expressed as:
    $U_{\text{circuit}_{1}}=\left(\bigotimes_{i=1}^{3}CNOT_{i,i+1}\right)\cdot\left(\bigotimes_{i=1}^{4}RX(\theta_{i})\right)\cdot\left(\bigotimes_{i=1}^{4}RY(x_{i})\right)$
    $U_{\text{circuit}_{1}}=(U_{\text{Entangle}}.U_{\text{Rotation}}.U_{{\text{Enc}}_{1}})$ (6)
    In case of two quanvolutional layers, the first quanvolutional layer (QCL1) has a depth of $4$, while that of second (QCL2) has a depth of $1$. This configuration is strategically chosen to distinctly evaluate the impact on performance when gradients from both layers are accessible in ResQuNNs, in contrast to situations where the gradients of only the second layer are accessible. It is noteworthy that at reduced and/or equal depths in both layers, the performance difference may not be as pronounced or readily observable even if the gradients in second layer are also available. Applied to the initial state $|0000\rangle$, the final state after the first QCL can be expressed as:
    |$\ket{\psi_{\text{QCL}_{1}}}=\prod_{l=1}^{4}U_{\text{circuit}_{1}}\ket{0000}$ (7)
    where $l$ represents the total number of layers, i.e., number of times the circuit is repeated before measurement. The size of output after every quanvloutional layers is detailed in Table I. |$\ket{\psi_{\text{QCL1}}}$ is the final state state after the first quanvolutionla layer before measurement and upon measurement the quantum state collapses and produce a classical result:
    $y=\mathcal{M}\ket{U_{\text{QCL1}}}$ (8)
    where $\mathcal{M}$ denotes the qubit measurement in computational basis. Since the result from first quanvolutional layer is classical, we need to encode again before passing it to the second quanvolutional layer using the same equation (Eq. 1):
    $U_{{\text{Enc}}_{2}}=|q_{i}\rangle=RY(y_{i})|0\rangle=\cos\left(\frac{y_{i}}{2}\right)|0\rangle+\sin\left(\frac{y_{i}}{2}\right)|1\rangle$ (9)
    The kernel size remains the same so a total of 4 input features will be encoded from in every step:
    $U_{{\text{Enc}}_{2}}=\bigotimes_{i=1}^{4}RY(y_{i})$ (10)
    The rest of the operations of trainable quantum circuit in QCL2 are same as above and the final circuit equation will be:
    $U_{\text{circuit}_{2}}=\left(\bigotimes_{i=1}^{3}CNOT_{i,i+1}\right)\cdot\left(\bigotimes_{i=1}^{4}RX(\theta_{i})\right)\cdot\left(\bigotimes_{i=1}^{4}RY(y_{i})\right)$
    $U_{\text{circuit}_{2}}=U_{\text{Entangle}}.U_{\text{Rotation}}.U_{{\text{Enc}}_{2}}$ (11)
    The depth of quantum circuit in QCL is set to $1$. Hence, the final quantum state after the application of QCL2 will be:
    |$\ket{\psi_{\text{QCL}_{2}}}=U_{\text{circuit}_{2}}\ket{y_{i}}$ (12)
    $Output_{\text{Quanv}}=\mathcal{M}\ket{U_{\text{QCL2}}}$ (13)
  4. 4.
    The measurement yields a list of classical expectation values. Similar to a classical convolution layer, each expectation value corresponds to a separate channel in a single output pixel.
  5. 5.
    The outputs from quanvolutional layer are then postprocessed individually using both classical layer and quantum circuit.
    Classical layer to postprocess quanvolutional layer’s output
    First, we use classical layer to postprocess the output of quanvolutional layer, as highlighted in label 2 in Fig. 5). This is a major practice in state-of-the-art QuNN architectures, where a fully connected neuron layer is used to take the data from the quanvolutional layer and produces the final classification output. The number of neurons in this layer are equal to the total number of classes in the dataset used, which in our case is $10$ since we used MNIST dataset which has $10$ classes (more details in Section III).
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We restrict our experimentation to kernel size of $2\times 2$ and stride of $2$, and output feature size in case of different residual configuration (label 5 in Fig. 5) is presented in the Table I.
    TABLE I: Output Feature Size for Two Quanvolutional Layer
    Input Size Residual Configuration Output Feature Size
    $28\times 28\times 1$ No residual $7\times 7\times 1$
    $28\times 28\times 1$ $X+O1$ $14\times 14\times 1$
    $28\times 28\times 1$ $O1+O2$ $14\times 14\times 1$
    $28\times 28\times 1$ $X+O2$ $28\times 28\times 1$
    $28\times 28\times 1$ $(X+O1)+O2$ $28\times 28\times 1$
    The classical neural network layer with $10$ neurons can be represented as a fully connected layer, where each neuron receives inputs from all $n$ measurements of the quanvolutional layer and processes these inputs to produce an output. The output of $j$-th neuron in the classical layer can be described by the following Equation:
    $y_{j}=f\left(\sum_{i=1}^{n}w_{ji}\cdot m_{i}+b_{j}\right)$
    where:
    • $f$ is the activation function (SoftMax in our case).
    • $w_{ji}$ is the weight from the $i$-th input to the $j$-th neuron.
    • $b_{j}$ is the bias term for the $j$-th neuron.
    • $y_{j}$ is the output of the $j$-th neuron.
    The operation of this layer can be compactly represented in matrix form as:
    $Y=f(W\cdot M+B)$
    where:
    • $Y$ is the output vector $[y_{1},y_{2},...,y_{10}]$.
    • $f$ is applied element-wise to the result.
    • $W$ is the weight matrix.
    • $M$ is the vector of measurements from the quantum layer.
    • $B$ is the bias vector.
    Quantum Circuit to postprocess quanvolutional layer’s output
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Fig. 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by reducing the influence of classical layers being used at the end.
    The use of a quantum circuit for post-processing introduces certain computational overheads, mainly due to the prerequisite of re-encoding the (classical) outputs from the quanvolutional layers, a process that entails considerable computational expense. Also the following quantum circuit after the encoding takes longer to produce the results than a simple classical neuron layer. We use $4$ classes from MNIST dataset for experimental simplification. For the purpose of encoding the outputs from the quanvolutional layer, Amplitude Encoding [20], is used which can be described by the following Equation:
    |$\ket{\psi_{x}}=\sum_{i=1}^{2^{n}}x_{i}\ket{i}\$ (14)
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table I, the smallest output feature size is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits. Angle encoding in this case would require $784$ qubits, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by the quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximal feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The encoded data is then subjected to the unitary evolution,which can be described by the following Equation:
    $U(\theta)=\bigotimes_{j=1}^{9}\left(CNOT_{j,j+1}\right).\bigotimes_{i=1}^{10}R_{y}(\theta)$ (15)
II-B Residual Quanvolutional Neural Networks (ResQuNNs).
To overcome the gradient accessibility issue encountered in multi-layered QuNNs, as discussed in Section I-A, we introduce ResQuNNs. Drawing inspiration from classical residual neural networks, ResQuNNs incorporate residual blocks between quanvolutional layers. The residual blocks combine the output of existing quanvolutional layers with their own or the previous layer’s input before forwarding it to the subsequent quanvolutional layer, as shown in label 4 in Fig. 5, and play a pivotal role in facilitating comprehensive gradient access and, in turn, improving the overall training performance of QuNNs. The integration of residual blocks ensures that gradients flow efficiently across the network, thereby overcoming the challenges associated with training multiple quanvolutional layers.
II-C Strategic Placement of Residual Blocks.
Within the ResQuNN architecture, not every random residual configuration results in comprehensive gradient flow throughout the network, and residual blocks needs to be strategically placed between quanvolutional layers. In our empirical investigation, we conduct extensive experimentation to determine the appropriate locations for inserting residual blocks within ResQuNNs comprising two quanvolutional layers, as shown in Fig. 5, labeled 5. This analysis helps us uncover the most effective residual configurations for ResQuNNs, enabling the gradients to flow throughout the proposed ResQuNN architecture, resulting in a deeper architecture and and eventually the enhanced training performance. Based on the results of two quanvolutional layers, only the most effective configurations, i.e., those facilitating gradient flow through both layers were chosen for subsequent examination within three-layer quanvolutional frameworks, as denoted by label 6 in Fig. 5.
We then trained ResQuNNs with two quanvolutional layers for all residual configurations for a multiclass classification problem (more on dataset in Section III) to assess the impact of gradient accessibility, both throughout the entire network and within segments of the network, on the learning efficacy of QuNNs.
TABLE II: Dataset and Hyperparemeters Specifications. BS = batch size, Opt = optimizer and LR = learning rate
Dataset Specs BS Opt LR Quanvolutional layer specs Postprocessing
Total classes Samples/class Train Data Test Data Encoding Kernel size # of Qubits Classical Quantum
10 200 1600 400 16 Adam 0.01 Angle (RY) $2\times 2$ 4 Dense layer (10 neurons) -
4 200 640 160 16 Adam 0.01 Angle (RY) $2\times 2$ 4 - 10-Qubit circuit (4 measured)
III Experimental Setup
We trained the proposed ResQuNNs on a selected subset of the MNIST dataset. Table II summarizes the dataset and hyperparameters specification used in this paper. The ResQuNNs are trained for $30$ epochs. Adam optimizer with learning rate of $0.01$ is used for the optimization. The objective function chosen for training is cross entropy loss, which can be described (for C classes) by the following equation:
$H(p,q)=-\sum_{c=1}^{C}p(c)\log(q(c))$
where $H(p,q)$ is the cross entropy loss, $p(c)$ represents the true probability distribution for class $c$ and $q(c)$ denotes the predicted probability distribution for class. All experiments are executed utilizing Pennylane, a Python library tailored for differentiable programming on quantum computing platforms [21]. An overview of the experimental toolflow utilized in this paper is depicted in Fig. 6.
Figure 6: Overview of Our Experimental Tool-flow
IV Results
IV-A Two Quanvolutional Layers
We use kernel size of $2\times 2$, which results in $4$-qubit quanvolutional layers. Our experimental framework involved extensive testing of all residual configurations as outlined in label 3 of Fig. 5. For each residual configuration, we investigated the gradient propagation through the quanvolutional layers. Afterwards, we trained the ResQuNNs for all these configurations to assess the impact of gradient accessibility throughout the network’s layers.
IV-A1 Analysis of Gradients Accessibility
In the case of No Residual setting, as previously shown in Fig. 3, we have already demonstrated the fact that the gradients are exclusively accessible for the second quanvolutional layer. Therefore, we now proceed to investigate the gradient propagation through the quanvolutional layers for all other residual settings (from label 5 in Fig. 5), the results of which are illustrated in Fig. 7. It is worth mentioning here that while the gradient accessibility results are presented with quantum layer postprocessing, the accessibility of gradients remains the same across layers with classical postprocessing as well.
Figure 7: Gradients Propagation Through the Network for Two Quanvolutional Layers for Different Residual Configurations. X=Input, O1, O2=Output of first and second quanvolutional layer respectively. To demonstrate gradients accessibility, we consider each quanvolutional layer with a quantum circuit depth of one
Our analysis reveals that specific residual configurations, notably $O1+O2$ and $(X+O1)+O2$, significantly enhance gradient propagation through the quanvolutional layers. This improvement not only makes them more conducive to gradient-based optimization strategies but also facilitates deeper learning within Quantum Neural Networks (QuNNs). In contrast, certain configurations, such as $X+O1$ and $X+O2$, mirror the conditions of a no residual setting, where gradients fail to propagate past the terminal quanvolutional layer. Consequently, these configurations may exhibit a constrained training potential, impacting their effectiveness in QuNNs.
Figure 8: Two Quanvolutional Layers - Training results of all residual configurations with classical postprocessing and $\textbf{2}\times\textbf{2}$ Kernel.
We observe that certain residual configurations, specifically $O1+O2$ and $(X+O1)+O2$, significantly enhance gradient propagation through the quanvolutional layers by facilitating the propagation of gradients through both the quanvolutional layers. This improvement not only makes them more conducive to gradient-based optimization strategies but also facilitates deeplearning within QuNNs. In contrast, other residual configurations, namely $X+O1$ and $X+O2$, are similar to the no residual setting, wherein the gradients fail to propagate beyond the last quanvolutional layer in the network. Consequently, these configurations may exhibit a constrained training potential.
IV-A2 Training Analysis
We then trained the proposed ResQuNNs on a subset of MNIST dataset, as detailed in Section III, for all the residual configurations labeled as 3 in Fig. 5, separately with classical and quantum postprocessing. The experimental procedures are consistent across all trials, adhering to the parameters discussed in Section III. It is worth mentioning here that the No residual configuration also serves as a common benchmark for all the other residual configurations to analyze whether or not the inclusion of residual connections bring any advantage. Below, we discuss the training results of ResQuNNs both with classical and quantum postprocessing.
Classical Postprocessing
We first evaluate the training performance of ResQuNNs that utilize classical neuron layers to postprocess the output of quanvolutional layers, as highlighted by label 2 in Fig. 5. The training results of ResQuNNs with classical postprocessing are shown in Fig. 8. A common and a key observation is that the incorporation of residual connections consistently improves the training performance of QuNNs, regardless of the specific residual configuration. For instance, while only the gradients of second quanvolutional layer are accessible in both No residual and $X+O1$ configurations, the introduction of a residual connection in $X+O1$ leads to superior performance compared to the No residual approach. We now individually explore the impact of different residual configurations on the performance of QuNNs with a focus on training efficiency and generalization capabilities (validation accuracy).
  • No Residual. When there is No residual block, the QuNNs exhibit no training as shown in label 1 of Fig. 8.
  • X+O1. When the residual block is added after the first qunavolutional layer, i.e., the input is added to the output of first quanvolutional and then passed as input to the second quanvolutional layer, the model shows significant performance enhancement with improved training and validation accuracy compared to No residual setting, as shown in label 2 of Fig. 8. While $X+O1$ residual configuration marks a significant improvement over the No residual configuration, it nonetheless presents a sub-optimal training performance.
  • O1+O2. Contrary to the No residual and $X+O1$ configurations, the residual setup of $O1+O2$ facilitates the propagation of gradients through both quanvolutional layers, which is considered more conducive for gradient-based optimization techniques. This facilitation of gradient propagation results in a marked improvement in performance with the $O1+O2$ configuration, as depicted in label 3 of Fig. 8 The training and generalization capabilities of the $O1+O2$ configuration significantly exceed those of $X+O1$, demonstrating approximately a 31% increase in training accuracy and a 45% enhancement in validation accuracy, respectively. This superiority can be attributed to the $O1+O2$ configuration’s facilitation of gradient propagation through both quanvolutional layers, in contrast to $X+O1$, where gradient access is limited exclusively to the final quanvolutional layer.
  • X+O2 and (X+O1)+O2. The residual configurations where the input is propagated directly ($X+O2$) or indirectly ($(X+O1)+O2$) to the output of second quanvolutional layer, the ResQuNNs outperform other residual configurations, as shown in and of Fig. 8. Additionally, both these residual settings exhibits very similar performance.
Figure 9: Two Quanvolutional Layers - Comparison of ResQuNNs with benchmark models for $\textbf{2}\times\textbf{2}$ Kernel.
We argue that the residual configurations facilitating gradient flow through both quanvolutional layers, specifically, $O1+O2$ and $(X+O1)+O2$, will inherently exhibit superior performance in terms of both training and generalization compared to those configurations where gradient access is confined to the last quanvolutional layer only (No Residual, $X+O1$, and $X+O2$ configurations). This is evident form the performance comparison of $X+O1$ and $O1+O2$ (label 2 and 3 in Fig. 8), where $O1+O2$ performs significantly better. Additionally, we hypothesize that the superior performance of $X+O2$ and $(X+O1)+O2$ can be attributed to the last classical layer (used to meaningfully post process quanvolutional layer’s output) in the network since the (input) classical data in both these settings is directly ($X+O2$) or indirectly ($(X+O1)+O2$) passed to the classical layer which tends to overshadow the learning contribution from quanvolutional layers.
Below, we scrutinize the influence of classical layer at the end in all these residual configurations to identify the residual configurations yielding better training and generalization performance where the quanvolutional layers primarily contribute to the learning.
Comparison with Benchmark Models
To assess the impact of the classical layer in various residual configurations of ResQuNNs, we conducted a detailed analysis using benchmark models. The benchmark models replicate the ResQuNNs with the residual configurations as shown in label 3 of Fig. 5, with one key difference: the quanvolutional layers were made untrainable, i.e., their weights are not updated during training, and only the classical layers were subjected to training. This exercise allows us to determine the extent to which the classical layer tends to dominate the learning process, compared to the quanvolutional layers across all configurations.
The benchmark models are trained using the same hyperparameters and dataset as detailed in Section III, and the results are compared with the models models, where both quanvolutional and classical layers were trainable. The comparison results are presented in Fig. 9. The observed performance trends in these models gives insights into the effectiveness of quanvolutional layers in different residual configurations. The key findings from this comparative analysis are summarized below:
No Residual. In the scenario where no residual connections are utilized, gradient flow is restricted exclusively to the second quanvolutional layer. Our observations indicate that the performance of the model, when both quanvolutional and classical layers undergo training, is very similar to that of the benchmark models in which training is restricted solely to the classical layers, with quanvolutional layers remaining static, as shown in label 1 of Fig. 9. Nevertheless, in both settings, this particular configuration results in minimal or No training.
Residual Configuration X+O1: In this residual configuration, despite gradient accessibility being limited to the second quanvolutional layer, the introduction of residual connections significantly enhances the model’s training capability. However, the comparative analysis of actual ResQuNNs (from Fig. 8) with benchmark models suggests that, the performance in this very residual configuration is predominantly influenced by the classical layer at the end of the quanvolutional layers, as highlighted by label 2 in Fig. 9. This observation suggests that, within this configuration, the classical layer plays a critical role in the model’s overall learning and performance outcomes.
Residual Configuration (O1+O2): This residual configuration potentially enables the deep learning by permitting gradient propagation through the quanvolutional layers, thus ensuring their active involvement in the learning process. This results in a significant enhancement in the model’s performance, particularly when compared with the benchmark model. As denoted by label 3 in Fig. 9, the performance of the model when both the quanvolutional layers along with classical layer are trained surpasses that of the benchmark model (where only classical layer is trained and quanvolutional layers are kept frozen) by approximately 66% in training accuracy and 64% in validation accuracy. This observation implies that in $O1+O2$ residual configuration, the learning performance is primarily driven by the quanvolutional layers. Such a notable enhancement underscores the benefit of fully utilizing both quanvolutional layers when gradients are accessible for them, highlighting the importance of their contribution to the learning process.
Figure 10: Two Quanvolutional Layers - Training results of all residual configurations with quantum postprocessing $\textbf{2}\times\textbf{2}$ Kernel.
Residual Configuration with Direct/Indirect Input to Classical Layer (X+O2 and (X+O1)+O2): In scenarios where the input is either directly or indirectly routed to the classical layer at the network’s end, such as in $X+O2$ and $(X+O1)+O2$, the performance of QuNNs with trainable quanvolutional demonstrates a close approximation to that of the benchmark models. This observation is highlighted by labels 4 and 5 in Fig. 9, indicating the predominant role of the classical layer in the learning process. The similar performance is achieved when only the classical layer is subjected to training, in comparison to scenarios where both quanvolutional and classical layers are trained, highlighting the significant role of the classical layer in the learning process in these residual configurations.
However, considering the enhanced performance in the $O1+O2$ configuration, which allows gradient flow through both the quanvolutional layers, it suggests that $(X+O1)+O2$ could be effective in scenarios without a classical layer at the end. This is because it also allows gradient propagation across all quanvolutional layers, potentially enabling the deep learning in QuNNs and eventually enhancing the learning performance, when the classical layer’s dominance is removed. To further elaborate on this, in the following section, we use the quantum circuit to postprocess the results of quanvolutional layers to further understand the role of residual approach in enabling deep learning in QuNNs, when there are no classical layers involved.
Quantum Postprocessing
In this section, we present the training results of ResQuNNs using quantum layer/circuit to postprocess the results of quanvolutional layer(s), as shown in label 3 of Fig. 5. This very architectural configuration, we observe clear advantage of residual configurations that allow the gradient flow across both the layers of the network compared to those which allow partial gradient flow. The training results are depicted in Fig. 10. Below, we separately discuss the results of ResQuNN with different residual configurations (from label 5 in Fig. 5), employing two quanvolutional layers.
  • Configurations with Restricted Gradient Flow (No residual, X+O1, X+O2): The residual configurations that allow partial gradient flow, i.e, No residual, X+O1 and X+O2 exhibits no to negligible training, as shown in labels 1, 2 and 4 of Fig. 10. This outcome underscores the critical role of gradient flow in the training process of QuNNs.
  • Configurations Facilitating Full Gradient Flow (X+O2, (X+O1)+O2): The residual configurations that facilitate gradients flow through all the layers of the network specifically ‘$X+O2$’ and ‘$(X+O1)+O2$’, demonstrated significantly enhanced performance and completely outperforms other residual configurations, as illustrated by label 3 and 5 in Fig. 10.
The contrast in training performance across different residual configurations elucidates the tremendous importance of gradient flow in the architecture of ResQuNNs. Configurations enabling full gradient availability across network layers emerged as distinctly advantageous, siginificantly improving the training potential of ResQuNNs. Conversely, configurations with limited gradient flow manifested negligible training outcomes, highlighting the inadequacy of partial gradient availability for effective network learning. This distinction underlines the significance of architectural choices in the design of ResQuNNs, particularly in the context of integrating quantum and classical layers for enhanced learning capabilities.
IV-B Three Quanvolutional Layers
In previous sections, we established the critical importance of gradient accessibility across all the layers in QuNNs for achieving better training performance. In this section, we extend our exploration to demonstrate that our proposed ResQuNNs can further facilitate the utilization of an increased number of quanvolutional layers for more deeper networks.
Specifically, we consider the scenario of training QuNNs with three quanvolutional layers. Due to space constraints, we only present the results for residual configurations where gradients are accessible across all three quanvolutional layers, as shown in Fig. 11. Out of $15$ potential residual configurations examined, only two, specifically $(O1+O2)+O3$ and $((X+O1)+O2)+O3$, allowed gradient access across all three layers. The other configurations restricted gradient flow to one or two layers. This further underscores the significance of our proposed residual approach, as it enables the desired comprehensive gradient accessibility, with greater number of layers, effectively enabling the deep learning in QuNNs, which can be pivotal for optimizing complex problems.
Figure 11: Gradients Propagation Through the Network for Three Quanvolutional Layers for (O1+O2)+O3 (top) and ((X+O1)+O2)+O3 (bottom) and . To demonstrate gradients accessibility, we consider each quanvolutional layer with a quantum circuit depth of one
V Conclusion
In this paper, we presented a novel framework for developing Quantum Convolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the innovative concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability.
A key contribution of this work is the development of trainable quanvolutional layers, which enhance the learning potential of QuNNs, and the strategic incorporation of residual blocks to ensure comprehensive gradient accessibility across the network in multi-layered QuNNs, thus, enabling deeper, more complex QuNNs architectures to be effectively trained. Our empirical findings highlight the importance of optimal placement of residual blocks, providing critical insights for future QuNN architecture designs with even more deeper layers.
In summary, our work signifies a substantial leap in quantum deep learning, offering new directions for theoretical and practical quantum computing research. By tackling core issues of adaptability and gradient flow, our work paves the way for the development of more robust and efficient quantum neural networks, marking a pivotal step toward harnessing the full potential of quantum technologies.
Acknowledgements
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008.